I was given three sets of coordinates and instructed to form the quadratic equation. The coordinates are: (-1,0),(-0.16,-2),(0.67,0).
I tried to enter the coordinates into y = ax^2+bx+c and solve for a, b and c but had trouble, can you please help?

Answer by vertciel(183)

Hello there,
Do not be frightened by the decimals.
Simply substitute each pair of coordinates into the equation:
y = ax^2 + bx + c
You should arrive at:
a + b + c = 0 ------------(1)
0.256a - 0.16b + c = -2 --(2)
0.4489a + 0.67b + c = 0 --(3)
Solve the system. You can plug in the values for a, b, c into the first equation to see if they are right.